# HRNet: Differentially Private Hierarchical and Multi-Resolution Network for Human Mobility Data Synthesization

### v1

@@ 5.1. Setup
5.1.1. dataset:
We utilize five datasets, broadly classified into three categories: human mobility, taxi trajectory (road network), and synthetic data. Human mobility encompasses a wide range of unrestricted mobility patterns, while road network data is constrained by road networks. The datasets employed are Geolife (Zheng et al., 2009) (human mobility), Peopleflow (Sekimoto et al., 2011) (human mobility), Didi in Chengdu (Wang et al., 2022) (taxi trajectory), and two synthetic datasets. The synthetic datasets are used in the ablation study and discussed in detail in Section 5.4. We use randomly sampled $10,000$ trajectories from each dataset as the training dataset. See details of the datasets for Appendix A.2.
@@ 5.1.3. Competitors
Table 1. The discrepancy of the generated dataset to the real dataset (Geolfie / Peopleflow) with $w=32$ when $\varepsilon=2$ and $\delta=10^{-5}$.
JS Divergence ARE
@@ 5.1.4. Evaluation metrics:
To compare the quality of generated mobility data, we evaluate the discrepancies between generated and original stay-point trajectories (denoted as $\mathbf{v}$) across various metrics. We employ the Jensen-Shannon (JS) divergence and average relative error (ARE) for this evaluation, with details provided in the Appendix A.3. The key metrics considered are waypoint, destination, transition, travel distance, route, diameter distribution, density at specific time, trajectory density, and transition pattern.
@@ 5.2. Comparison with Baseline, PrivTrace, and Clustering
In this section, we present a comparative evaluation of our proposed method against existing approaches for human mobility not constrained by road networks. The competitors in this analysis include the Baseline, PrivTrace, and Clustering methods. The used datasets are the Geolife dataset and the Peopleflow dataset.
@@ 5.2.2. Utility-privacy trade-off
This subsection delves into the utility-privacy trade-off of our method in comparison to other competitors, achieved by varying the privacy budget. The results depicted in Figures 6 and 6 illustrate this analysis. It is observed that at lower privacy budgets (e.g., $\varepsilon<1$), PrivTrace and clustering sometimes exhibit better performance. This is attributed to their direct computation of transitions.
In contrast, deep learning methods generally require extensive training epochs to learn the hidden features of human mobility. However, our method enhances the learning process through pre-training and multi-task learning, effectively reducing the need for a larger privacy budget, unlike the baseline method. Furthermore, as the learning progresses, deep learning approaches increasingly outperform their counterparts. This advantage stems from their direct learning of raw data, avoiding the utility loss associated with sensitivity bounding, a common issue in statistical-based methods like clustering and PrivTrace.
@@ 5.3. Comparison with MTNet
Our results, as illustrated in Figures 6 and 6, focus on evaluating destination and route discrepancies. For both metrics, MTNet initially exhibits superior performance, primarily because it generates trajectories using road network metadata. However, as training progresses, our method demonstrates improved results. This discrepancy arises from the differing objectives of the training targets. Our method is tailored to learn semantically significant next waypoints. Conversely, MTNet is designed to directly learn the next road segment. As a result, MTNet is required to learn longer sequence which is difficult to learn in the noisy setting. Hence, our method has better convergence in the case where DP-SGD is applied.
@@ 5.4. Ablation study
In this section, we conduct an ablation study to elucidate the individual contributions of the key components of our proposed method: the hierarchical location encoding via deconvolutional network, multi-resolution multi-task learning, and private pre-training. Our analysis begins with the Geolife dataset, followed by an examination using synthetic datasets, each with distinct characteristics.
@@ 5.4.1. Real data
Figure 7. The discrepancy of transition on Geolife ($w=64$). $p$, $d$, and $m$ represent pre-raining, deconvolutional network, and multi-task learning.
Figure 8. The discrepancy of the transition on the Random dataset ($w=32$) for each training epoch. d, m, and p mean deconvolutional network, multi-task learning, and pre-training, respectively.
Our evaluation encompasses all six settings mentioned previously, focusing on their performance on the Geolife dataset with a grid size of $w=64$. The outcomes of this evaluation are depicted in Figure 8. The results indicate that the incorporation of each component – the deconvolutional network, multi-task learning, and pre-training – not only enhances the accuracy in terms of discrepancy reduction but also accelerates convergence.
@@ 5.4.3. Multi-task learning
As demonstrated in Section 5.4.1, the implementation of multi-task learning has significantly improved both empirical utility and convergence speed. This enhancement is largely due to the efficient utilization of inferences derived from simpler tasks, which align with the first law of geography. That is, the bias based on the first law of geography helps with the privacy-utility trade-off, but it is important to note that bias can also have a negative impact.
To further substantiate this point, we investigate a scenario where the first law of geography is not applicable, utilizing the Random dataset for this purpose. The results of this investigation are presented in Figure 8. Contrary to previous observations, the addition of multi-task learning in this context results in a decrease in utility. This decline can be attributed to the inductive bias towards the first law of geography, which, in the case of the Random dataset, leads to a misalignment with the actual data characteristics. Consequently, the multi-task learning framework, while beneficial under certain conditions, may inadvertently impact the results negatively when the underlying data does not adhere to the geographic principles it assumes.
@@ 5.5.2. Trajectory classification
Trajectory classification involves training a model to categorize trajectories into different types. Here, we consider classification of transportation modes, a well-known problem in trajectory classification (Zheng et al., 2010; Xiao et al., 2017; Liu et al., 2019). Some of trajectories in the Geolife dataset are labeled as bike, walk, car, bus, and train. We create $5,980$ trajectories using grid-based discretization with $w=32$, splitting them into a training set and a test set with a $0.9$ to $0.1$ ratio. Using the training set, we first train a generative model (HRNet or the baseline) to synthesize trajectories. Here, the label information is incorporated into the generative model by extending Equation 2 as follows:
${\rm encode}(v_{i})=[M_{\rm time}[t_{i}],M_{\rm POI}[l_{i}],M_{\rm label}[s]],$
where $M_{\rm label}$ is the embedding matrix for labels and $s$ represents the label index. This encoding method generates trajectories considering the label information. With this generated labeled data, we extract features following the method proposed by Liao et al. (Liao et al., 2007) and train a random forest classifier using these features. The lower part of Table 3 shows the classification results on the test dataset. The Random Forest results represent the model trained with raw data. Compared to the baseline, our model performs better, but there is still a significant performance drop. This is because trajectory classification requires maintaining spatio-temporal correlations, making it more challenging than next location prediction. Additionally, our model uses discretized time information with the embedding matrix, and the preprocessing required for embedding inherently leads to some information loss.
@@ 6.1. Non-deep Learning Models
One major drawback of non-deep learning models is that they require preprocessing to reduce sensitivity. This preprocessing often results in significant information loss. Moreover, due to their complexity, these models often cause errors in their privacy proofs (Buchholz et al., 2024).
@@
Clustering-based methods (Hua et al., 2015; Li et al., 2017) first cluster the locations to reduce the number of locations as a preprocessing step. Then, they count the transitions in the smaller location domain (i.e., classes) to create a probabilistic model with the Laplace mechanism. Many clustering-based methods (Hua et al., 2015; Li et al., 2017) have been found to cause unexpected privacy leakage in the data-dependent preprocessing due to flaws in the privacy proof (Miranda-Pascual et al., 2023). Instead, data-dependent preprocessing and the use of the Laplace mechanism, which is very effective with respect to privacy-utility trade-off, can sometimes perform well with small $\varepsilon$ values (see Figure 6). However, preprocessing leads to substantial information loss, and even with large $\varepsilon$ values, accuracy cannot be significantly improved. Our method, based on gradient descent (DP-SGD), does not require preprocessing, thereby mitigating these issues.
Tree and markov
Tree (Chen et al., 2012; He et al., 2015) and Markov (Wang et al., 2023; Gursoy et al., 2018) methods first construct a data structure (e.g., prefix tree) for trajectories. Then, they count elements (e.g., nodes) to create the probabilistic model. Essentially, tree and Markov methods face two problems due to the high uniqueness of patterns in trajectory: how to define locations (e.g., determining the grid size $w$) and how to resolve sparsity. Since the initial study by Chen et al. (Chen et al., 2012), subsequent researches has attempted to address these issues. In the subsequent work (Chen et al., 2013), substrings are considered for building an exploration tree based on a Markov assumption, resulting in higher leaf counts and better utility. DPT (He et al., 2015) addresses the sparsity problem by restricting movement to adjacent cells using a hierarchical structure, effectively capturing short trajectory features. However, this method’s limitation to adjacent cell movement prevents conversion to stay point trajectories, resulting in longer sequence lengths. Given DP’s characteristics, performance significantly degrades with longer sequence lengths, making it unsuitable for capturing broad patterns such as daily human movement. PrivTrace (Wang et al., 2023) addresses the sparsity problem heuristically, yielding good results on some datasets, but there remains the problem of determining constants used in heuristics. Our method addresses them using hierarchical networks and multi-task learning. As discussed in this paper, our method’s dependence on $w$ is low, and using the necessary fine-grained $w$ avoids the first issue. Multi-task learning includes low resolutions, thus mitigating the impact of sparsity without the adjacent assumption like DPT.
@@ 6.2. Deep Learning Models
As argued in this study, when using DP-SGD, sparsity emerges as a significant issue in deep learning, and thus not many studies propose using DP-SGD. Ahuja et al. (Ahuja et al., 2020) propose using negative sampling and skip-gram to solve this sparsity issue. Note that such techniques could also be applicable to our study. Wang et al. (Wang et al., 2022) address this problem similarly to DPT by restricting movement to adjacent road network segments. Hence, as demonstrated in Figure 6, they effectively capture short trajectory features and perform well with small $\varepsilon$ due to prior road network information. However, the same problem as DPT arises, making it unsuitable for capturing broad daily human movement patterns. Without the formal guarantees like DP, many deep learning approaches (Fontana et al., 2023; Ozeki et al., 2023; Song et al., 2023) rely on obfuscation caused by the generation itself. However, without any formal guarantee, it is unclear if privacy is genuinely preserved (Stadler et al., 2022). Even without privacy-aware considerations, sparsity remains a problem, and there are studies proposing models similar to ours. Lim et al. (Lim et al., 2022) address sparsity by employing multi-task learning with multiple resolutions, but their naive creation of embedding matrices for each resolution results in a large number of parameters ($O(n^{2})$), which is not suitable to DP-SGD. Lian et al. (Lian et al., 2020) adopt a hierarchical location encoding using a quadtree similar to our study, but instead of using a deconvolutional network, they treat quadtree paths as sequences and perform sequence embedding with the attention architecture (Vaswani et al., 2017). Consequently, encoding a single location requires processing a sequence length corresponding to the quadtree depth. Longer sequence lengths degrade DP-SGD performance, which is not ideal. Moreover, the applicability of multi-task learning to such methods is not trivial. Many deep learning methods for capturing trajectory features have been considered (Feng et al., 2018; Luo et al., 2021; Lim et al., 2020; Yang et al., 2020). However, directly encoding the tuple of latitude and longitude is fundamentally challenging (Lian et al., 2020), leading to the use of embedding matrices. Using embedding matrices poses the same issues as the baseline, and applying DP-SGD to these studies is expected to degrade accuracy. To resolve these issues, our study’s hierarchical location encoding component, along with associated multi-task learning and pre-training, is considered applicable.
@@ A.1. Derivation of the privacy budget allocation
We derive the solution (9) as follows. The noise introduced by the Laplace mechanism is $O\left(\frac{\log(w^{2})}{\varepsilon_{2}}\right)$  (Vadhan, 2017). Due to the definition of the transition matrix (see Equation 3), the signal in an element of the DP transition matrix is $O\left(\frac{|D|}{4^{i_{\rm res}}w^{2}}\right).$ Then, based on the noise and signal, SNR is represented as ${\rm SNR}=O\left(\frac{4^{i_{\rm res}}w^{2}\log(w^{2})}{|D|\varepsilon_{2}}\right).$ We can see that the asymptotic trend of SNR is determined solely based on the meta-information. Hence, setting $\varepsilon_{2}=c\left(\frac{4^{i_{\rm res}}w^{2}\log(w^{2})}{|D|}\right)$ ensures a constant SNR, with empirical results indicating that $c=0.018$ offers a balanced approach. While heuristic, experimental evidence suggests that this allocation strategy works well across various settings.
@@ A.3.1. Discrepancy
density at $t$: For density at $t$, the support encompasses all POIs ($L$). We assign $p_{t}$ and $q_{t}$ to represent the probability distributions from the original and generated datasets for time $t$, respectively. Then, the discrepancy is as follows: $\frac{1}{T}\sum_{t\in T}{\rm JS}(p_{t}||q_{t}),$ where $T$ is the set of all discrete times.

### v2

@@ 5.1. Setup
5.1.1. datasets:
We utilize five datasets, broadly classified into three categories: human mobility, taxi trajectory (road network), and synthetic data. Human mobility encompasses a wide range of unrestricted mobility patterns, while road network data is constrained by road networks. The datasets employed are Geolife (Zheng et al., 2009) (human mobility), Peopleflow (Sekimoto et al., 2011) (human mobility), Didi in Chengdu (Wang et al., 2022) (taxi trajectory), and two synthetic datasets. The synthetic datasets are used in the ablation study and discussed in detail in Section 5.4. We use randomly sampled $10,000$ trajectories from each dataset as the training dataset. See details of the datasets in Appendix A.2.
@@ 5.1.3. Competitors
Table 1. The discrepancy of the generated dataset to the real dataset (Geolife / Peopleflow) with $w=32$ when $\varepsilon=2$ and $\delta=10^{-5}$.
JS Divergence $\downarrow$ ARE $\downarrow$
@@ 5.1.4. Evaluation metrics:
To compare the quality of generated mobility data, we evaluate the discrepancies between generated and original stay-point trajectories (denoted as $\mathbf{v}$) across various metrics. We employ the Jensen-Shannon (JS) divergence and average relative error (ARE) for measuring discrepancy (both the lower the better), with details provided in the Appendix A.3. The key metrics considered are as follows which are all probability distributions. Waypoint captures which waypoints are traversed given a starting point. Destination captures which location serves as the final destination given a starting point. Transition captures where the trajectory moves next from a starting point. Travel distance represents the total length of a trajectory. Diameter measures the maximum distance between any two points along a trajectory. Route captures the route taken from a starting point. Density at $t$ is the probability of a trajectory being at a specific location at time $t$. Trajectory density is the probability of a trajectory passing through a specific area. Trajectory pattern captures the most frequent transition patterns.
@@ 5.2. Comparison with Baseline, PrivTrace, and Clustering
In this section, we present a comparative evaluation of HRNet against existing approaches for human mobility not constrained by road networks. The competitors in this analysis include the Baseline, PrivTrace, and Clustering methods. The used datasets are the Geolife dataset and the Peopleflow dataset.
@@ 5.2.2. Utility-privacy trade-off
This subsection delves into the utility-privacy trade-off of HRNet in comparison to other competitors, achieved by varying the privacy budget. The results depicted in Figures 6 and 6 illustrate this analysis. It is observed that at lower privacy budgets (e.g., $\varepsilon<1$), PrivTrace and clustering sometimes exhibit better performance. This is attributed to their direct computation of transitions.
In contrast, deep learning methods generally require extensive training epochs to learn the hidden features of human mobility. However, our method enhances the learning process through pre-training and multi-task learning, effectively reducing the need for a larger privacy budget, unlike the baseline method. Furthermore, as epsilon increases, deep learning approaches increasingly outperform their counterparts. This advantage stems from their direct learning of raw data, avoiding the utility loss associated with sensitivity bounding, a common issue in statistical-based methods like clustering and PrivTrace.
@@ 5.3. Comparison with MTNet
Our results, as illustrated in Figures 6 and 6, focus on evaluating destination and route discrepancies. For both metrics, MTNet initially exhibits superior performance, primarily because it generates trajectories using road network metadata. However, as epsilon increases, our method demonstrates improved results. This discrepancy arises from the differing objectives of the training targets. Our method is tailored to learn semantically significant next waypoints. Conversely, MTNet is designed to directly learn the next road segment. As a result, MTNet is required to learn a longer sequence which is difficult to learn under DP constraints. Hence, our method has better convergence in the case where DP-SGD is applied.
@@ 5.4. Ablation study
In this section, we conduct an ablation study to elucidate the individual contributions of the key components of HRNet: the hierarchical location encoding via deconvolutional network, multi-resolution multi-task learning, and private pre-training. Our analysis begins with the Geolife dataset, followed by an examination using synthetic datasets, each with distinct characteristics.
@@ 5.4.1. Real data
Figure 7. The discrepancy of transition on Geolife ($w=64$). $p$, $d$, and $m$ represent pre-raining, deconvolutional network, and multi-task learning, respectively.
Figure 8. The discrepancy of transition on Random dataset ($w=32$). $d$, $m$, and $p$ mean deconvolutional network, multi-task learning, and pre-training, respectively.
The evaluation on the Geolife dataset with a grid size of $w=64$ is depicted in Figure 8. The results indicate that the incorporation of each component – the deconvolutional network, multi-task learning, and pre-training – not only enhances the accuracy in terms of discrepancy reduction but also accelerates convergence.
@@ 5.4.3. Multi-task learning
HRNet benefits from the efficient utilization of inferences derived from simpler tasks, which align with the first law of geography. It is important to note that bias based on the first law of geography can also have a negative impact. To further substantiate this point, we investigate a scenario where the first law of geography is not applicable, utilizing the Random dataset for this purpose. The results of this investigation are presented in Figure 8. Contrary to previous observations, the addition of multi-task learning in this context results in a decrease in utility. This decline can be attributed to the inductive bias towards the first law of geography, which, in the case of the Random dataset, leads to a misalignment with the actual data characteristics. Consequently, the multi-task learning framework, while beneficial under certain conditions, may impact the results negatively when the underlying data does not adhere to the geographic principles it assumes.
@@ 5.5.2. Trajectory classification
Trajectory classification involves training a model to categorize trajectories into different types. Here, we consider classification of transportation modes, a well-known problem in trajectory classification (Zheng et al., 2010; Xiao et al., 2017; Liu et al., 2019). Some of trajectories in the Geolife dataset are labeled as bike, walk, car, bus, and train. We create $5,980$ trajectories using grid-based discretization with $w=32$, splitting them into a training set and a test set with a $0.9$ to $0.1$ ratio. Using the training set, we first train a generative model (HRNet or the baseline) to synthesize trajectories. Here, the label information is incorporated into the generative model by extending Equation 2 as follows: ${\rm encode}(v_{i})=[M_{\rm time}[t_{i}],M_{\rm POI}[l_{i}],M_{\rm label}[s]],$ where $M_{\rm label}$ is the embedding matrix for labels and $s$ represents the label index. This encoding method generates trajectories considering the label information. With this generated labeled data, we extract features following the method proposed by Liao et al. (Liao et al., 2007) and train a random forest classifier using these features. The lower part of Table 3 shows the classification results on the test dataset. The Random Forest results represent the model trained with raw data. Compared to the baseline, our model performs better, but there is still a significant performance drop. This is because trajectory classification requires maintaining spatio-temporal correlations, making it more challenging than next location prediction. Additionally, our model uses discretized time information with the embedding matrix, and the preprocessing required for embedding inherently leads to some information loss.
@@ 6.1. Non-deep Learning Models
Clustering-based methods (Hua et al., 2015; Li et al., 2017) first cluster the locations to reduce the number of locations as a preprocessing step. Then, they count the transitions in the smaller location domain (i.e., classes) to create a probabilistic model with the Laplace mechanism. Many clustering-based methods (Hua et al., 2015; Li et al., 2017) have been found to cause unexpected privacy leakage in the data-dependent preprocessing due to flaws in the privacy proof (Miranda-Pascual et al., 2023). Instead, data-dependent preprocessing and the use of the Laplace mechanism, which is very effective with respect to privacy-utility trade-off, can sometimes perform well with small $\varepsilon$ values (see Figure 6). However, clustering based preprocessing leads to substantial information loss, and even with large $\varepsilon$ values, accuracy cannot be significantly improved.
Tree and Markov models
Tree (Chen et al., 2012; He et al., 2015) and Markov (Wang et al., 2023; Gursoy et al., 2018) methods first construct a data structure (e.g., prefix tree) for trajectories. Then, they count elements (e.g., nodes) to create the probabilistic model. Essentially, tree and Markov methods face two problems due to the high uniqueness of patterns in trajectory: how to define locations (e.g., determining the grid size $w$) and how to resolve sparsity. Since the initial study by Chen et al. (Chen et al., 2012), subsequent studies have attempted to address these issues. In the work (Chen et al., 2013), substrings are considered for building an exploration tree based on a Markov assumption, resulting in higher leaf counts and better utility. DPT (He et al., 2015) addresses the sparsity problem by restricting movement to adjacent cells using a hierarchical structure, effectively capturing short trajectory features. However, this method’s limitation to adjacent cell movement prevents conversion to stay point trajectories, resulting in longer sequence lengths. Given DP’s characteristics, performance significantly degrades with longer sequence lengths, making it unsuitable for capturing broad patterns such as daily human movement. PrivTrace (Wang et al., 2023) addresses the sparsity problem heuristically, yielding good results on some datasets, but there remains the problem of determining constants used in heuristics. Our method addresses them using hierarchical networks and multi-task learning.
@@ 6.2. Deep Learning Models
As argued in this study, when using DP-SGD, sparsity emerges as a significant issue in deep learning, and thus not many studies propose using DP-SGD. Ahuja et al. (Ahuja et al., 2020) propose using negative sampling and skip-gram to solve this sparsity issue. Note that such techniques could also be applicable to our study. Wang et al. (Wang et al., 2022) address this problem similarly to DPT by restricting movement to adjacent road network segments. Hence, as demonstrated in Figure 6, they effectively capture short trajectory features and perform well with small $\varepsilon$ due to prior road network information. However, the same problem as DPT arises, making it unsuitable for capturing broad daily human movement patterns. Without the formal guarantees like DP, many deep learning approaches (Fontana et al., 2023; Ozeki et al., 2023; Song et al., 2023) rely on obfuscation caused by the generation itself. However, without any formal guarantee, it is unclear if privacy is genuinely preserved (Stadler et al., 2022). Even without privacy-aware considerations, sparsity remains a problem, and there are studies proposing models similar to ours. Lim et al. (Lim et al., 2022) address sparsity by employing multi-task learning with multiple resolutions, but their naive creation of embedding matrices for each resolution results in a large number of parameters ($O(n^{2})$), which is not suitable to DP-SGD. Lian et al. (Lian et al., 2020) adopt a hierarchical location encoding using a quadtree similar to our study, but instead of using a deconvolutional network, they treat quadtree paths as sequences and perform sequence embedding with the attention architecture (Vaswani et al., 2017). Consequently, encoding a single location requires processing a sequence length corresponding to the quadtree depth. Longer sequence lengths degrade DP-SGD performance, which is not ideal. Moreover, the application of multi-task learning to such methods is not trivial. Many deep learning methods for capturing trajectory features have been considered (Feng et al., 2018; Luo et al., 2021; Lim et al., 2020; Yang et al., 2020). However, directly encoding the tuple of latitude and longitude is fundamentally challenging (Lian et al., 2020), leading to the use of embedding matrices. Using embedding matrices poses the same issues as the baseline.
@@ 7. Conclusion
Acknowledgements.
This work was supported by the support of JST SICORP JPMJSC2107, JST CREST JPMJCR21M2, JST PRESTO JPMJPR23P5, JSPS KAKENHI JP22H03595, JP23K24851, JP21K19767, NSF CNS-2125530, CNS-2124104, IIS-2302968, and CDC 1NU38FT000001-01-00.
@@ A.1. Derivation of the privacy budget allocation
We derive the solution (9) as follows. The noise introduced by the Laplace mechanism is $O\left(\frac{\log(w^{2})}{\varepsilon_{2}}\right)$  (vadhan2017complexity). Due to the definition of the transition matrix (see Equation 3), the signal in an element of the DP transition matrix is $O\left(\frac{|D|}{4^{i_{\rm res}}w^{2}}\right).$ Then, based on the noise and signal, SNR is represented as ${\rm SNR}=O\left(\frac{4^{i_{\rm res}}w^{2}\log(w^{2})}{|D|\varepsilon_{2}}\right).$ We can see that the asymptotic trend of SNR is determined solely based on the meta-information. Hence, setting $\varepsilon_{2}=c\left(\frac{4^{i_{\rm res}}w^{2}\log(w^{2})}{|D|}\right)$ ensures a constant SNR, with empirical results indicating that $c=0.018$ offers a balanced approach. While heuristic, experimental evidence suggests that this allocation strategy works well across various settings.
@@ A.3.1. Discrepancy
density at $t$: For density at $t$, the support encompasses all POIs ($L$). We assign $p_{t}$ and $q_{t}$ to represent the probability distributions from the original and generated datasets for time $t$, respectively. Then, the discrepancy is as follows: $\frac{1}{|T|}\sum_{t\in T}{\rm JS}(p_{t}||q_{t}),$ where $T$ is the set of all discrete times.